The difference of three times a number and half the number is 60. What is the number form of this problem.
Answer by Timnewman(323)

Hi!
Let the number be "x"
Three times the number=3x
Half of the number = x/2.
From the first sentence,
The difference between three times a nember and half of the number is 60,
i.e,
3x-(x/2)=60
multiply out by 2
6x-x=120
5x=120
divide bothsides by 5
5x/5=120/5
x=24
The number is 24
HOPE THIS HELPS;...
------ --TIMNEWMAN------